The lengths of pregnancies are normally distributed with a mean of 266 days and a standard deviation of 15 days.
a. Find the probability of a pregnancy lasting 309 days or longer.
b. If the length of pregnancy is in the lowest 4​%, then the baby is premature. Find the length that separates premature babies from those who are not premature

Answer by Boreal(15235)

z=(x-mean)/sd
z=(309-266)/15=2.867
probability z > 2.867 is 0.0021.
-
the lowest 4% from the calculator or the table is -1.751
-1.751=(x-266)/15
-26.26=x-266
x=239.74 days or about 240 days is the cutoff.